Felix has an average of 90 in five tests, each test with 100 points. What is the lowest possible score felix could have gotten in a test?

Answer by ikleyn(52787)

.
Since Felix has the average of 90 in five tests, the sum of his scores in 5 tests is 5*90 = 450 points.


The lowest score in one of the test is achieved when all the sums for each of all other tests are equal to 100, and then


the lowest possible point = 450 - 4*100 = 50.

Solved.